# A Digit-Sum Formula for Böttcher Coordinates

### 02 Apr 2026 v1

@@ Abstract.
Let $p$ be an odd prime and let $\phi(x)=x^{p^{2}}+p^{2}x^{p^{2}+1}$. We prove an unconditional digit-sum congruence formula for the coefficients of the Böttcher coordinate of $\phi$. The argument separates the divisible case $k\equiv 0\pmod{p}$ from the general case, gives a complete vector-partition description of the surviving $B$-term monomials, and reduces the remaining partition sum to a one-variable cumulant. As consequences we recover the residue-class congruences conjectured by Salerno and Silverman and obtain an explicit description of the first block of coefficients modulo $p$.
@@ 1. Introduction
The local study of holomorphic iteration begins with the functional equations of Schröder and Koenigs, and the theory of complex dynamics took its modern shape through the early twentieth-century work of Böttcher, Fatou, and Julia [1, 6]. For a one-dimensional analytic germ with a superattracting fixed point, Böttcher’s theorem provides a canonical local coordinate in which the map becomes a pure monomial. In polynomial dynamics this coordinate at infinity does considerably more than linearize the map: it identifies the basin of infinity with the exterior disk, encodes the escape-rate function, and underlies external rays, landing theory, and the combinatorial study of Julia sets and the Mandelbrot set [2, 3, 6]. The same normal form also appears in several variables, where the local geometry near superattracting fixed points becomes subtler [5].
In recent years Böttcher coordinates have also become important in non-archimedean dynamics. Salerno and Silverman studied their integrality properties for one-dimensional superattracting germs over $p$-adic rings [8], while Fu and Nie later settled a conjecture of Salerno–Silverman on the radius of convergence in the wild superattracting case [4]. The family relevant for the present paper is
$\phi(x)=x^{p^{2}}+p^{2}x^{p^{2}+1}.$
If
$f_{\phi}(x)=x\sum_{k\geq 0}\frac{a_{k}}{k!}x^{k}$
@@
$\phi\bigl(f_{\phi}(x)\bigr)=f_{\phi}(x^{p^{2}}),$
then Salerno and Silverman recorded striking mod-$p$ patterns in the coefficients $a_{k}$; their original formulation is Conjecture 25 of [8]. For an integer $n=\sum_{i\geq 0}n_{i}p^{i}$ with $0\leq n_{i}\leq p-1$, let
$S_{p}(n):=\sum_{i\geq 0}n_{i}$
be the sum of its base-$p$ digits.
From now on, we assume that $p\geq 3$. We note that the corresponding result for $p=2$ can be established similarly, but since the arguments for all $p\geq 3$ fit into a unified framework, we do not treat the case $p=2$ in the present paper. The following are our main theorems.
Theorem 1.1 (Digit-sum formula).
@@
In fact, we have a stronger result
(1.2) $a_{k}\equiv(-1)^{S_{p}(k)}(a+1)^{S_{p}(k)-1}\pmod{p}$.
We also recover the original residue-class conjecture of Salerno and Silverman.
Theorem 1.2 (Salerno–Silverman’s Conjecture 25).
@@
The proof is organized by separating the residue class $a=0$ from the classes $a\neq 0$. Section 2 recalls the AABBCC recursion and proves the basic $A$- and $C$-term lemmas. Section 3 supplies a complete proof of the divisible case $a=0$. The remaining sections treat $a\neq 0$: Section 4 gives a vector-partition expansion of the $B$-term, Section 5 evaluates the first block of coefficients and collapses the multivariate cumulant, and Section 6 proves (1.1). In the final section we prove (1.2) and deduce Theorem 1.2.
@@ 2. The recursion and two structural lemmas
Here $P[x^{n}]$ denotes the coefficient of $x^{n}$ in the power series $P(x)$. Throughout we use Legendre’s formula
@@ 5. Low coefficients and a multivariate cumulant collapse for $a\neq 0$
By Lagrange inversion, the unique formal solution of $U=xe^{-U}$ satisfies
@@
7. Proof of (1.2) and Theorem 1.2
We first promote Proposition 5.1 to the fully explicit form of Theorem $1.$

### 06 Apr 2026 v2

# Digit-sum Formulas and Divisible Classes for Böttcher Coordinates of $x^{p^{2}}+p^{r+2}x^{p^{2}+1}$

@@ Abstract.
Let $p$ be an odd prime and, for $r\geq 0$, let
$\phi_{r}(x)=x^{p^{2}}+p^{r+2}x^{p^{2}+1}.$
For $r=0$ we prove a complete digit-sum congruence formula modulo $p$ for the coefficients of the Böttcher coordinate of $\phi_{0}$. As consequences we prove Conjecture 25 of Salerno and Silverman and obtain an explicit description of the first block of coefficients modulo $p$. For $r\geq 1$ we prove on the divisible non-pure class a leading term theorem together with a global digit-weight lower bound. These results yield an unconditional proof of Conjecture 27(a) of Salerno and Silverman.
@@ 1. Introduction
The local study of superattracting germs goes back to Böttcher’s theorem, which provides a canonical local coordinate in which a germ becomes a pure monomial. In complex dynamics this coordinate plays a central role both near a superattracting fixed point and, for polynomials, at infinity, where it encodes the escape-rate function and the external combinatorics of Julia sets and the Mandelbrot set; see, for example, [1, 2, 3, 6, 7]. The same normal form also appears in higher-dimensional holomorphic dynamics [5].
In the non-archimedean setting, Böttcher coordinates have an additional arithmetic aspect. Salerno and Silverman studied their integrality properties for one-dimensional superattracting germs and formulated several precise conjectures in the wild case [8]. More recently, Fu and Nie proved the expected radius of convergence in a general wild superattracting setting [4]; in particular, their theorem settles Conjecture 27(b) for the family considered here. The family relevant for the present paper is
$\phi_{r}(x)=x^{p^{2}}+p^{r+2}x^{p^{2}+1}\qquad(r\geq 0),$
where $p\geq 3$ is an odd prime. If
$f_{r}(x)=x\sum_{k\geq 0}\frac{a_{k}(r)}{k!}x^{k}$
@@
$\phi_{r}\bigl(f_{r}(x)\bigr)=f_{r}(x^{p^{2}}),$
then Conjectures 25 and 27 of [8] predict striking congruence and valuation patterns for the coefficients $a_{k}(r)$. Since Conjecture 27(b) is already supplied by Fu and Nie, the higher-fiber part of the present paper is concerned only with Conjecture 27(a).
The special fiber $r=0$ already exhibits a rich mod-$p$ structure. In that case one expects a digit-sum formula for every coefficient, not only on the divisible class $p\mid k$ but on every residue class modulo $p$. Our first result establishes the full congruence law.
Theorem 1.1 (Digit-sum formula in the special fiber).
@@
In fact, one has the stronger closed formula
(1.2) $a_{k}\equiv(-1)^{a+s}(a+1)^{a+s-1}\pmod{p}$.
As an immediate consequence we recover the residue-class congruences conjectured in [8, Conjecture 25].
Theorem 1.2 (Salerno–Silverman’s Conjecture 25 for $r=0$).
@@
For the higher fibers $r\geq 1$, the relevant structure is valuation-theoretic. If
$v_{i}(r):=\operatorname{ord}_{p}\bigl(a_{p^{i}}(r)\bigr)\qquad(i\geq 0),$
and if
$m=m_{0}+m_{1}p+\cdots+m_{s}p^{s},\qquad 0\leq m_{i}\leq p-1,$
we set
$\Lambda_{m}(r):=\sum_{i=0}^{s}m_{i}v_{i+1}(r),\qquad\epsilon_{m}:=m_{0}+\cdots+m_{s}=S_{p}(m).$
Our third main theorem gives the precise leading term on the divisible non-pure class and, from it, proves the valuation statement Conjecture 27(a) of [8].
Theorem 1.3 (Divisible classes and Conjecture 27(a) for $r\geq 1$).
Fix $r\geq 1$. Then the following hold.
  1. (a)
    If $m\geq 1$ is not a power of $p$, then
    $a_{pm}(r)=(-1)^{\epsilon_{m}}p^{\Lambda_{m}(r)}+O\bigl(p^{\Lambda_{m}(r)+1}\bigr),$
    where
    $v_{0}(r)=r,\qquad v_{1}(r)=pr,$
    and, for every $n\geq 2$,
    $v_{n}(r)=\begin{cases}(p+1)p^{n-2}-2+v_{n-2}(r),&\text{if $n$ is even and $r\geq n-1$,}\\[3.0pt] p\,v_{n-1}(r),&\text{otherwise.}\end{cases}$
    In particular,
    $\operatorname{ord}_{p}\bigl(a_{pm}(r)\bigr)=\Lambda_{m}(r).$
  2. (b)
    If $p\mid k$, then
    $\operatorname{ord}_{p}\bigl(a_{k}(r)\bigr)=\frac{1-p^{-r}}{p-1}k+O(1).$
The special-fiber part of the paper is driven by a complete analysis of the AABBCC recursion modulo $p$, together with a vector-partition description of the surviving $B$-term monomials. The higher-fiber part is organized around a simultaneous induction on three main assertions—a global lower bound, a leading theorem on the divisible non-pure class, and a pure-power leading theorem—together with the resulting subadditivity of the digit weight. Once these ingredients are in place, the pure-power branch pattern and Conjecture 27(a) follow.
We do not treat the case $p=2$ here. The arguments for odd $p$ already exhibit the full mechanism, and keeping $p\geq 3$ avoids a number of low-characteristic notational distractions.
@@ 2. The recursion and two structural lemmas
Here $P[x^{n}]$ denotes the coefficient of $x^{n}$ in the power series $P(x)$. For a nonnegative integer
$N=\sum_{i\geq 0}N_{i}p^{i},\qquad 0\leq N_{i}\leq p-1,$
we write
$S_{p}(N):=\sum_{i\geq 0}N_{i},\qquad T_{p}(N):=\frac{N!}{p^{\operatorname{ord}_{p}(N!)}},$
so $T_{p}(N)$ is the prime-to-$p$ part of $N!$. Throughout we use Legendre’s formula
@@ 5. Low coefficients and a multivariate cumulant collapse for $a\neq 0$
By Lagrange inversion (see [9, 10]), the unique formal solution of $U=xe^{-U}$ satisfies
@@
Proof of (1.2) and Theorem 1.2
We first combine Proposition 5.1 with (1.1) to obtain the closed formula in Theorem $1.$